# LagCeresa.bib

### v1

@@ 6.4. Image of the determinantal forms and tautological tropical cycles
$\kappa^{\prime}:H^{\bullet}(F;\mathbb{Z})\hskip 2.84526pt\longrightarrow$ $\hskip 5.69054ptC^{\bullet}(F;\mathbb{Z})$
@@
where $p_{i}$ is the cotangent coordinate associated to $x_{i}$ in $X(B)$. In particular, composing with the the isomorphism $\Lambda^{k}T_{\mathbb{Z}}^{*}B\cong H^{n-k}(F;\mathbb{Z})$, elements of the form
@@
Notice that in each term appearing in Definition 5.3, the $n-k$ indices $\{j_{k+1},\dots,j_{n}\}$ would correspond to the union of the $j$ indices $\{s_{1},\dots,s_{j}\}$ and the $n-(j+k)$ (ordered) indices
$\{i_{j+k+1},\dots,i_{n}\}:=\{1,\dots,n\}\backslash\{i_{1},\dots,i_{j+k}\}\cong\{1,\dots,n\}\backslash\{j_{1},\dots,j_{k},s_{1},\dots,s_{j}\}.$
@@
$\Omega_{1}^{3}:=e_{1}^{*}\otimes dx_{2}\wedge dx_{3}-e_{2}^{*}\otimes dx_{1}\wedge dx_{2}+e_{3}^{*}\otimes dx_{1}\wedge dx_{2}$
@@
$\kappa:H_{\bullet}(F;\mathbb{Z})\hskip 2.84526pt\longrightarrow$ $\hskip 5.69054ptC_{\bullet}(F;\mathbb{Z})$
@@
Consider a $1$-cell $e$ in $B$ with $\partial e=b_{1}-b_{2}$; this defines an element in $C_{1}(B;\mathbb{Z})$ whose image through $\hat{\phi}$ is an $(n+1)$-cycle $\hat{\phi}(e)\cong F\times[0,1]$ with boundary $F_{b_{1}}-F_{b_{2}}$. In this situation, the appropriate determinantal form is
$\Omega^{n}_{0}=\sum_{i\in\{1,\dots,n\}}dx_{i},$
@@
where $vol_{n-1}^{i}=dp_{1}\wedge\dots\wedge dp_{i-1}\wedge dp_{i+1}\wedge\dots\wedge dp_{n}$. Therefore, if $v=(v_{1},\dots,v_{n})$ is a vector in $TB$ tangent to $e$, and because such a vector has rational slope since $e$ is tropical, we can decompose (although not canonically) the base $B\cong T^{n}$ as $T^{n-1}\times S^{1}$, with $v$ tangent to the $S^{1}$ direction, whose coordinate we denote by $\theta$. Then the tropical pairing $\int_{e}\Omega^{n}_{0}$ corresponds precisely to $\int_{(T^{n-1}\times S^{1})\times[0,1]}vol_{n-1}\wedge d\theta\wedge dt,$ with $t$ being the coordinate along $[0,1]$ and $vol_{n-1}$ a volume form on the $T^{n-1}$-component; this matches the symplectic flux used in Proposition 2.6.
@@
From Remark 5.11, we know that tropical flux does not obstruct algebraic equivalence for 00-cycles. It does, however, yield symplectic flux obstructing Lagrangian cobordisms between the corresponding fibres. In the case of Remark 5.11, the corresponding symplectic flux between $F_{p}$ and $F_{q}$ in $X(S^{1})\cong T^{2}$ is
@@
where $C$ is a $2$-cylinder with $\partial C=F_{p}-F_{q}$. It implies that $F_{p}$ and $F_{q}$ are not Lagrangian cobordant unless $p=q$.
@@ 7. The Lagrangian Ceresa cycle
$AJ_{p}(C)-(-1)_{*}AJ_{p}(C)\in CH_{1}(J(C)),$
and when there is no ambiguity from the choice of $p$ we often abuse notation and simply write $C-C^{-}$. If they exist, we denote by $L_{C}$ or simply $L$ the Lagrangian lift of $C$, and by $L_{C}^{-}$ or simply $L^{-}$ the Lagrangian lift of $C^{-}$. We then define the Lagrangian Ceresa cycle as the Lagrangian $L-L^{-}$ inside $X(B)$. As mentioned in the introduction, for $C$ hyperelliptic, choosing $p$ to be a Weierstrass point implies $AJ_{p}(C)=(-1)_{*}AJ_{p}(C)$, therefore $L_{C}=L_{C}^{-}$. This means that the simplest non-trivial case to study the Lagrangian Ceresa is a non-hyperelliptic curve $C$ with $g(C)=3$, and here we restrict ourselves to the type $K4$ curve from Figure 5.
@@
7.1. The $3$-manifold $L-L^{-}$
From now on, when referring to the Lagrangian Ceresa cycle we will always be working with a curve of type $K4$, therefore $L-L^{-}$ is a $3$-manifold in a $6$-torus $X(J(C))\cong T^{6}$. Because such a curve is trivalent, we know from constructions of Mikhalkin and Matessi [matessi_lag_pair_of_pants, mikhalkin_trop_to_Lag_corresp] that it admits a Lagrangian lift $L$.
@@ 7.1. The $3$-manifold $L-L^{-}$
Figure 6. A representation of a part of the Lagrangian lift of $C$ inside $X(B)\cong B\times T^{3}$. Each of the pairs of pants is contained in a $2$-plane inside $B$ indicated on the figure, and $S^{1}_{(p_{1},p_{2},p_{3})}$ denotes the $S^{1}$-factor in the fibre direction generated by the line $(p_{1},p_{2},p_{3})$ in $(T^{*}B)_{b}$. Similarly, $T^{2}_{(p_{1},p_{2},p_{3}),(q_{1},q_{2},q_{3})}$ denotes the $2$-torus in a fibre generated by the linear span of the vectors $(p_{1},p_{2},p_{3})$ and $(q_{1},q_{2},q_{3})$ in $(T^{*}B)_{b}$.
The Lagrangian lift $L^{-}$ of $C^{-}$ is simply the image of $L$ by a global symplectomorphism. Indeed, $C^{-}$ is the image of $C$ by the automorphism $(-1)$ of $B$. Because $(-1)$ is tropical, the induced symplectomorphism on $T^{*}J(C)$ yields a symplectomorphism $\varphi^{-}$ of $T^{*}J(C)/T^{*}_{\mathbb{Z}}J(C)$. Notice $L$ and $L^{-}$ are not disjoint, therefore the Lagrangian Ceresa cycle is only immersed.
Although we know that the Ceresa cycle is nullhomologous 4.6, it may not yet be clear that the Lagrangian Ceresa cycle is. This fact will follow from our proof of Theorem 7.1 (see Remark 7.3), although we can already exhibit some supporting evidence. Tropically, $(-1)$ acts as $-\operatorname{Id}$ on $H_{1}(J(C);\mathbb{Z})$ and $T_{\mathbb{Z}}^{*}J(C)\cong H_{2}(F;\mathbb{Z})$, therefore it acts trivially on $H_{1}(J(C))\otimes H_{2}(F;\mathbb{Z})$. We know that this homology group contains $\hat{\phi}(\overline{C})$, and it will follow from the proof of Theorem 7.1 that it also contains $L$.
@@ 7.2. Main result
For a generic tropical curve $C$ of type $K4$ with $g(C)=3$, the Lagrangian Ceresa cycle $L-L^{-}\in algCob^{or}(X(J(C)))_{hom}$ has infinite order.
@@
For a generic tropical curve $C$ of type $K4$ with $g(C)=3$, the Lagrangian Ceresa cycle $L-L^{-}\in\mathcal{C}ob^{or}(X(J(C)))_{hom}$ has infinite order.
This result builds on Zharkov’s proof of the corresponding tropical statement 4.5 about the Ceresa cycle $C-C^{-}\in CH_{1}(J(C))$ being non-torsion. There Zharkov constructs a tropical $2$-chain $\gamma_{0}$ in $C_{2}(J(C),T_{\mathbb{Z}}B)$ satisfying $\partial\gamma_{0}=\overline{C}-\overline{C^{-}}$. He shows that
$\int_{\gamma_{0}}\Omega^{3}_{1}=-ad,$
@@
Recall from Example 6.4 that the determinantal form $\Omega^{3}_{1}$ is mapped through the appropriate choice of $\hat{\psi}$ to
$\frac{1}{2}\omega^{2}\in\Omega^{2}(X(J(C))),$
@@
$\int_{\gamma_{0}}\Omega^{3}_{1}=\frac{1}{2}\int_{\hat{\phi}(\gamma_{0})}\omega^{2}.$
@@
To prove Theorem 7.1, we apply Lemma 2.4 with $\alpha:=\frac{1}{2}\omega$, integrating $\frac{1}{2}\omega^{2}$ over an $n+1$-chain between $L$ and $L^{-}$. Notice that $\hat{\phi}(\gamma_{0})$ is an $(n+1)$-chain between $\hat{\phi}(\overline{C})$ and $\hat{\phi}(\overline{C^{-}})$, on which $\frac{1}{2}\omega^{2}$ integrates to something which is not in its periods. In Matessi’s construction of the Lagrangian pair-of-pants, he shows (Corollary 3.19 in [matessi_lag_pair_of_pants]) that these come in families which converge to his piecewise-linear Lagrangian lift. By inspection of [matessi_lag_pair_of_pants, Section 2] and by Section 6.4, the piecewise-linear Lagrangian lift of $C$ coincides with $\hat{\phi}(C)$. This family gives an $(n+1)$ chain $\gamma$ connecting $L$ and $\hat{\phi}(C)$, on which $\omega^{2}$ necessarily vanishes (as on any $(n+1)$-chain given by a $1$-dimensional family of Lagrangians). The same argument yields an $(n+1)$-chain $\gamma^{-}$ connecting $L^{-}$ and $\hat{\phi}(C^{-})$ on which $\omega^{2}$ vanishes. Therefore
$\int_{\gamma\cup\hat{\phi}(\gamma_{0})\cup\gamma^{-}}\frac{1}{2}\omega^{2}=\int_{\hat{\phi}(\gamma_{0})}\frac{1}{2}\omega^{2}=\int_{\gamma_{0}}\Omega_{1}^{3}=-ad.$
Because the maps $\phi$ and $\psi$ are isomorphisms, the periods of $\frac{1}{2}\omega^{2}$ are exactly the periods of $\Omega^{3}_{1}$. Therefore for a generic choice of edge lengths, $-ad$ is not torsion modulo the periods of $\frac{1}{2}\omega^{2}$, and we conclude. ∎
@@
While $\phi$ being an isomorphism already implied that $\hat{\phi}(C)$ and $\hat{\phi}(C^{-})$ are homologous, the $(n+1)$-chains provided by Matessi’s converging family of pairs-of-pants allows us to conclude that $L$ and $L^{-}$ are homologous in $X(J(C))$.
@@
While for generic choices of edge lengths, the flux $-ad$ is not torsion in $\mathcal{P}^{3}_{1}$, our computation 5.5 shows that choosing for instance $f=0$ immediately yields $-ad\in\mathcal{P}^{3}_{1}$. Although no analogous results have been brought to light tropically, classically [Beauville_torsion_Ceresa_class, Beauville_torsion_Ceresa_cycle] exhibit an example of a non-hyperelliptic curve which nevertheless has torsion Ceresa cycle.
@@
It is worth noting that a few years after Ceresa proved the inifinite order statement for the Ceresa cycle in $\text{Griff}_{1}(J(C))$, Bardelli proved that this group is in fact infinitely generated. He proceeds by constructing generators as follows: for every $k\geq 1$, the cycles $C_{k}-C_{k}^{-}\in\text{Griff}_{1}(J(C))$ are pushforwards of Ceresa cycles of curves $C_{k}$ for which $J(C_{k})$ has a $k$-to-$1$ map to $J(C)$. It is expected that one could prove a similar statement symplectically, only with our methods this would require a tropical construction of these infinitely many generators, which to the best of our knowledge has not appeared in the literature. There has been, however, progress on the tropical Schottky problem; for instance an algorithm to produce a tropical curve $C_{k}$ from the data of its Jacobian (and a theta divisor) is suggested in [bolognese2017curvestropicaljacobians, Section 7], and in [Chua2017SchottkyAC].

### v2

@@ 5.4. Image of the determinantal forms and tautological tropical cycles
$\kappa^{\prime}:H^{\bullet}(F_{b};\mathbb{Z})\hskip 2.84526pt\longrightarrow$ $\hskip 5.69054ptC^{\bullet}(F_{b};\mathbb{Z})$
@@
where $p_{i}$ is the cotangent coordinate associated to $x_{i}$ in $X(B)$. In particular, composing with the isomorphism $\Lambda^{k}(T_{\mathbb{Z}}^{*}B)_{b}\cong H^{n-k}(F_{b};\mathbb{Z})$, elements of the form
@@
Notice that in each term appearing in Definition 2.10, the nkn-k indices $\{j_{k+1},\dots,j_{n}\}$ would correspond to the union of the $j$ indices $\{s_{1},\dots,s_{j}\}$ and the n(j+k)n-(j+k) (ordered) indices
[…]
@@
$\Omega_{2,1}^{3}:=e_{1}^{*}\otimes dx_{2}\wedge dx_{3}-e_{2}^{*}\otimes dx_{1}\wedge dx_{2}+e_{3}^{*}\otimes dx_{1}\wedge dx_{2}$
@@
$\kappa:H_{\bullet}(F_{b};\mathbb{Z})\hskip 2.84526pt\longrightarrow$ $\hskip 5.69054ptC_{\bullet}(F_{b};\mathbb{Z})$
@@
Consider a $1$-cell $e$ in $B$ with $\partial e=b_{1}-b_{2}$; this defines an element in $C_{1}(B;\mathbb{Z})$ whose image through $\hat{\phi}$ is an $(n+1)$-cycle $\hat{\phi}(e)\cong T^{n}\times[0,1]$ with boundary $F_{b_{1}}-F_{b_{2}}$. In this situation, the appropriate determinantal form is
$\Omega^{n}_{1,0}=\sum_{i\in\{1,\dots,n\}}dx_{i},$
@@
where $vol_{n-1}^{i}=dp_{1}\wedge\dots\wedge dp_{i-1}\wedge dp_{i+1}\wedge\dots\wedge dp_{n}$. Because $e$ is tropical, a vector $v=(v_{1},\dots,v_{n})$ in $TB$ tangent to $e$ has rational slope. Therefore, we can (non-canonically) decompose the base $B\cong T^{n}$ as $T^{n-1}\times S^{1}$ in such a way that $v$ tangent to the $S^{1}$-direction. We denote by $\theta$ the coordinate on this $S^{1}$-factor. Then the tropical pairing $\int_{e}\Omega^{n}_{1,0}$ corresponds precisely to $\int_{(T^{n-1}\times S^{1})\times[0,1]}vol_{n-1}\wedge d\theta\wedge dt,$ with $t$ being the coordinate along $[0,1]$ and $vol_{n-1}$ a volume form on the $T^{n-1}$-component; this matches the symplectic flux used in Proposition 3.6.
@@
From Remark 2.18, we know that tropical flux does not obstruct algebraic equivalence for 0-cycles. It does, however, yield symplectic flux obstructing Lagrangian cobordisms between the corresponding fibres. In the case of Remark 2.18, the corresponding symplectic flux between $F_{p}$ and $F_{q}$ in $X(S^{1})\cong T^{2}$ is
@@
where $C$ is a $2$-cylinder with $\partial C=F_{p}-F_{q}$. It implies that $F_{p}$ and $F_{q}$ are not oriented Lagrangian cobordant unless $p=q$.
@@ 6. The Lagrangian Ceresa cycle
$AJ_{p}(C)-(-1)_{*}AJ_{p}(C)\in\text{Griff}_{1}(J(C)),$
and when there is no ambiguity from the choice of $p$ we often abuse notation and simply write $C-C^{-}$. If they exist, we denote by $L_{C}$ the Lagrangian lift of $C$, and by $L_{C}^{-}$ the Lagrangian lift of $C^{-}$. We then define the Lagrangian Ceresa cycle as the Lagrangian $L_{C}-L_{C}^{-}$ inside $X(B)$. As mentioned in the introduction, for $C$ hyperelliptic, choosing $p$ to be a Weierstrass point implies $AJ_{p}(C)=(-1)_{*}AJ_{p}(C)$, therefore $L_{C}=L_{C}^{-}$. This means that the simplest non-trivial case to study the Lagrangian Ceresa is a non-hyperelliptic curve $C$ with $g(C)=3$, and here we restrict ourselves to the type $K4$ curve from Figure 6.
@@
6.1. The polarised symplectic torus $X(J(C))$
Generally, given any tropical torux $B=B(Q)$ for some matrix $Q\in GL(n;\mathbb{R})$, the periods of the symplectic form on $X(B):=T^{*}B/T^{*}_{\mathbb{Z}}B$ are given by the entries of the matrix $Q$. To see this, consider global coordinates $\{x_{1},x_{2},x_{3},p_{1},p_{2},p_{3}\}$ on $X(B)$, where the $p_{i}$ are cotangent coordinates associated to the $x_{i}$. The nine non-vanishing periods of $\omega$ correspond to integrals over generators of
$H_{1}(B;\mathbb{Z})\otimes H_{1}(F_{b};\mathbb{Z})\subset H_{2}(X(B);\mathbb{Z}),$
where $F_{b}$ is a torus fibre at any point $b\in B$, and we have implicitly used the isomorphism $X(B)\cong B\times F_{b}$. It follows from the construction of $B(Q)$ that generators of $H_{1}(F_{b};\mathbb{Z})$ can be identified with unit vectors $\{e_{1},e_{2},e_{3}\}$ in $T^{*}_{\mathbb{Z}}B\cong\mathbb{R}^{3}$, while generators of $H_{1}(B;\mathbb{Z})$ are identified with $\{Q\cdot e_{1},Q\cdot e_{2},Q\cdot e_{3}\}=\{\gamma_{1},\gamma_{2},\gamma_{3}\}$. The integral of $\omega$ over the generator $Q\cdot e_{i}\otimes e_{j}$ of $H_{1}(B;\mathbb{Z})\otimes H_{1}(F_{b};\mathbb{Z})$ is thus the symplectic area of the surface in $X(J(C))$ tangent to the vectors $(Q\cdot e_{i},0,0,0)$ and $(0,0,0,e_{j})$, which is simply $(Q\cdot e_{i})_{j}$.
In the particular case where $B=J(C)$ is the Jacobian of a tropical curve, Remark 2.14 tells us that the matrix $Q$ is the polarisation matrix of the curve. In particular, it is symmetric and positive-definite. For the $K4$ curve from Figure 6, the periods of the symplectic form on $X(J(C))$ are therefore integral combinations of the six different entries of the polarisation matrix
$Q=\begin{pmatrix}a+e+f&-f&-e\\ -f&b+d+f&-d\\ -e&-d&c+d+e\end{pmatrix}$.
Explicitly, in terms of the dual basis $\{\gamma_{1}^{*},\gamma_{2}^{*},\gamma_{3}^{*}\}$ of $H^{1}(B;\mathbb{Z})$ and $\{e_{1}^{*},e_{2}^{*},e_{3}^{*}\}$ of $H^{1}(F_{b};\mathbb{Z})$, the class of the symplectic form is
$[\omega]=$ $(a+e+f)\gamma_{1}^{*}\otimes e_{1}^{*}+(b+d+f)\gamma_{2}^{*}\otimes e_{2}^{*}+(c+d+e)\gamma_{3}^{*}\otimes e_{3}^{*}$
$-f(\gamma_{2}^{*}\otimes e_{1}^{*}+\gamma_{1}^{*}+e_{2}^{*})-e(\gamma_{3}^{*}\otimes e_{1}^{*}+\gamma_{1}^{*}\otimes e_{3}^{*})-d(\gamma_{2}^{*}\otimes e_{3}^{*}+\gamma_{3}^{*}\otimes e_{2}^{*})\in H^{2}(X(B);\mathbb{R}).$
6.2. The $3$-manifold $L_{C}-L_{C}^{-}$
Because the $K4$ curve $C$ is trivalent, we know from constructions of Mikhalkin and Matessi [matessi_lag_pair_of_pants, mikhalkin_trop_to_Lag_corresp] that it admits a Lagrangian lift $L_{C}$ to $X(J(C))$. Therefore $L_{C}-L_{C}^{-}$ exists, and is a Lagrangian $3$-manifold inside a symplectic $6$-torus $X(J(C))$.
@@ 6.2. The $3$-manifold $L_{C}-L_{C}^{-}$
Figure 7. A representation of a part of the Lagrangian lift of $C$ inside $X(B)\cong B\times T^{3}$. Each of the pairs of pants is associated to a trivalent vertex contained in a $2$-plane inside $B$ indicated on the figure. $S^{1}_{(p_{1},p_{2},p_{3})}$ denotes the $S^{1}$-factor in the fibre direction generated by the line $(p_{1},p_{2},p_{3})$ in $(T^{*}B)_{b}$. Similarly, $T^{2}_{(p_{1},p_{2},p_{3}),(q_{1},q_{2},q_{3})}$ denotes the 22-torus in a fibre generated by the linear span of the vectors $(p_{1},p_{2},p_{3})$ and $(q_{1},q_{2},q_{3})$ in $(T^{*}B)_{b}$.
The homology of $L_{C}$ can be computed by iteratively applying Mayer-Vietoris using the cover represented in Figure 7, from which one finds $b_{1}(L_{C})=b_{2}(L_{C})=6$ and $H_{3}(L_{C})\cong H_{0}(L_{C})\cong\mathbb{Z}$. In particular, $L_{C}$ is orientable.
The Lagrangian lift $L_{C}^{-}$ of $C_{C}^{-}$ is simply the image of $L_{C}$ by a global symplectomorphism. Indeed, $C^{-}$ is the image of $C$ by the automorphism (1)(-1) of $B$. Because (1)(-1) is tropical, the induced symplectomorphism on $T^{*}J(C)$ yields a symplectomorphism $\varphi^{-}$ of $T^{*}J(C)/T^{*}_{\mathbb{Z}}J(C)$. Notice $L_{C}$ and $L_{C}^{-}$ are not disjoint, therefore the Lagrangian Ceresa cycle is only immersed.
Although we know the Ceresa cycle is nullhomologous (see Remark 2.6), it may not yet be clear that the Lagrangian Ceresa cycle is. This fact will follow from our proof of Theorem 6.1 (see Remark 6.3), nevertheless we can already exhibit some supporting evidence. Tropically, (1)(-1) acts as $\operatorname{Id}$ on $H_{1}(J(C);\mathbb{Z})$ and $T_{\mathbb{Z}}^{*}J(C)\cong H_{2}(F;\mathbb{Z})$, therefore it acts trivially on $H_{1}(J(C))\otimes H_{2}(F;\mathbb{Z})$. We know this homology group contains $\hat{\phi}(\overline{C})$, and it will follow from the proof of Theorem 6.1 that it also contains $L_{C}$.
@@ 6.3. Main result
For a generic tropical curve $C$ of type $K4$ with $g(C)=3$, the Lagrangian Ceresa cycle $L_{C}-L_{C}^{-}\in algCob^{or}(X(J(C)))_{hom}$ has infinite order.
@@
For a generic tropical curve $C$ of type $K4$ with $g(C)=3$, the Lagrangian Ceresa cycle $L_{C}-L_{C}^{-}\in\mathcal{C}ob^{or}(X(J(C)))_{hom}$ has infinite order.
This result builds on Zharkov’s proof of the corresponding statement about the tropical Ceresa cycle $C-C^{-}\in\text{Griff}_{1}(J(C))$ being non-torsion (Theorem 2.5). There, Zharkov constructs a tropical $2$-chain $\gamma_{0}$ in $C_{2}(J(C),T_{\mathbb{Z}}J(C))$ satisfying $\partial\gamma_{0}=\overline{C}-\overline{C^{-}}$. He shows that
$\int_{\gamma_{0}}\Omega^{3}_{2,1}=-ad,$
@@
To be more precise, he constructs a $2$-chain $\gamma_{0}^{\prime}$ in $C_{2}(J(C);T_{\mathbb{Z}}J(C))$ between $AJ_{p}(C)$ and $AJ_{p^{\prime}}(C)^{-}$ in $J(C)$, where $p$ and $p^{\prime}$ are chosen so that the images of the edge $C$ in $AJ_{p}(C)$ and $AJ_{p^{\prime}}(C)^{-}$ coincide. For instance, take $p=p_{1}$ and $p^{\prime}=p_{3}$ (see Figure 6). Notice that $AJ_{p}(C)$ and $AJ_{p^{\prime}}(C)$ are algebraically equivalent for any $p$ and $p^{\prime}$ in $C$: this is realised by the equivalence
$W:=\{p,AJ_{p}(C)\}\subset C\times J(C),$
which one verifies is a tropical cycle.
With this in mind, when we refer to the tropical chain $\gamma_{0}$, we will really mean $\tilde{W}\cup\gamma^{\prime}_{0}$, where $\tilde{W}\in C_{2}(J(C);T_{\mathbb{Z}}J(C))$ is the tropical chain constructed in 2.2.2 from $W$, which satisfies
$\int_{\tilde{W}}\Omega_{0}=0$
by Lemma 2.16. Therefore
$\int_{\gamma_{0}:=\gamma^{\prime}_{0}\cup\tilde{W}}\Omega_{0}=\int_{\gamma^{\prime}_{0}}\Omega_{0}.$
The tropical chain $\gamma^{\prime}_{0}\in C_{2}(J(C);T_{\mathbb{Z}}J(C))$ is supported on the following $5$ parallelograms in $J(C)$:
  • $E_{1}$ with edges $AJ_{p}(A)$, $AJ_{p}(B)$, $(-1)_{*}AJ_{p^{\prime}}(A)$, and $(-1)_{*}AJ_{p^{\prime}}(B)$;
  • $E_{2}$ with edges $(-1)_{*}AJ_{p^{\prime}}(B)$, $(-1)_{*}AJ_{p^{\prime}}(F)$, $(-1)AJ_{p^{\prime}}(B)$ translated by $(-1)_{*}(AJ_{p^{\prime}}(p_{0})-AJ_{p^{\prime}}(p_{2}))$, and $(-1)_{*}AJ_{p^{\prime}}(F)$ translated by $(-1)_{*}(AJ_{p^{\prime}}(p_{2})-AJ_{p^{\prime}}(p_{3}))$;
  • $E_{3}$ with edges $AJ_{p}(B)$, $(-1)_{*}AJ_{p^{\prime}}(E)$, $AJ_{p}(B)$ translated by $(-1)_{*}(AJ_{p^{\prime}}(p_{0})-AJ_{p^{\prime}}(p_{1}))$, and $(-1)_{*}AJ_{p^{\prime}}(E)$ translated by $AJ_{p}(p_{2})-AJ_{p}(p_{3})$.
  • $E_{4}$ with edges $AJ_{p}(E)$, $AJ_{p}(F)$, $AJ_{p}(E)$ translated by $AJ_{p}(p_{2})-AJ_{p}(p_{0})$, and $AJ_{p}(F)$ translated by $AJ_{p}(p_{1})-AJ_{p}(p_{0})$.
  • $E_{5}$ with edges $AJ_{p}(E)$, $(-1)_{*}AJ_{p^{\prime}}(E)$, $AJ_{p}(D)$, and $(-1)_{*}AJ_{p^{\prime}}(D)$.
These are parallelograms in $J(C)$ because we have chosen $p$ and $p^{\prime}$ so that the image of $C$ by $AJ_{p}$ and $(-1)_{*}AJ_{p^{\prime}}$ coincide, therefore cancel out in $C-C^{-}$. Furthermore, there is a choice of orientation for these parallelepipeds which we do not describe here, but is found in Zharkov’s original construction [zharkov_tropical_ceresa]. The $T_{\mathbb{Z}}J(C)$-coefficients are given as follows, in the basis $\{e_{1},e_{2},e_{3}\}$: (1,1,0)(-1,1,0) for $E_{2}$, $E_{3}$ and $E_{4}$, (1,0,0)(-1,0,0) for $E_{1}$, and (0,1,1)(0,1,-1) for $E_{5}$. One readily verifies (given the orientations) that $\partial\gamma^{\prime}_{0}=\overline{AJ_{p}(C)}-\overline{(-1)_{*}AJ_{p^{\prime}}(C)}$. Notice furthermore that the $T_{\mathbb{Z}}J(C)$-coefficients are tangent to the parallelograms for $E_{1}$, $E_{2}$, $E_{4}$ and $E_{5}$. Hence Proposition 2.15 implies that
$\int_{\gamma^{\prime}_{0}}\Omega_{0}=\int_{E_{3}}\Omega_{0}=-ad,$
where the minus sign follows from the orientation.
Recall from Example 5.5 that the determinantal form $\Omega^{3}_{2,1}$ is mapped through the appropriate choice of $\hat{\psi}$ to
$\frac{1}{2}\omega^{2}\in\Omega^{4}(X(J(C))),$
@@
$\int_{\gamma_{0}}\Omega^{3}_{2,1}=\frac{1}{2}\int_{\hat{\phi}(\gamma_{0})}\omega^{2}.$
@@
To prove Theorem 6.1, we apply Lemma 3.4 with $\alpha:=\frac{1}{2}\omega$, integrating $\frac{1}{2}\omega^{2}$ over an $4$-chain between $L_{C}$ and $L_{C}^{-}$. Notice that $\hat{\phi}(\gamma_{0})$ is a $4$-chain between $\hat{\phi}(\overline{C})$ and $\hat{\phi}(\overline{C^{-}})$, on which $\frac{1}{2}\omega^{2}$ integrates to something which is not in its periods. In Matessi’s construction of the Lagrangian pair-of-pants, he shows (Corollary 3.19 in [matessi_lag_pair_of_pants]) that they come in families which converge to his piecewise-linear Lagrangian lift. By inspection of [matessi_lag_pair_of_pants, Section 2] and by Section 5.5, the piecewise-linear Lagrangian lift of $C$ coincides with $\hat{\phi}(C)$. This family gives an $4$ chain $\gamma$ connecting $L_{C}$ and $\hat{\phi}(C)$, on which $\omega^{2}$ necessarily vanishes (as on any $4$-chain given by a $1$-dimensional family of Lagrangians). The same argument yields an $4$-chain $\gamma^{-}$ connecting $L_{C}^{-}$ and $\hat{\phi}(C^{-})$ on which $\omega^{2}$ vanishes. Therefore
$\int_{\gamma\cup\hat{\phi}(\gamma_{0})\cup\gamma^{-}}\frac{1}{2}\omega^{2}=\int_{\hat{\phi}(\gamma_{0})}\frac{1}{2}\omega^{2}=\int_{\gamma_{0}}\Omega_{2,1}^{3}=-ad.$
Because the maps $\phi$ and $\psi$ are isomorphisms, the periods of $\frac{1}{2}\omega^{2}$ are exactly the periods of $\Omega^{3}_{2,1}$. Therefore for a generic choice of edge lengths, ad-ad is not torsion modulo the periods of $\frac{1}{2}\omega^{2}$, and we conclude. ∎
@@
While $\phi$ being an isomorphism already implied that $\hat{\phi}(C)$ and $\hat{\phi}(C^{-})$ are homologous, the $(n+1)$-chains provided by Matessi’s converging family of pairs-of-pants allows us to conclude that $L_{C}$ and $L_{C}^{-}$ are homologous in $X(J(C))$.
@@
While for generic choices of edge lengths, the flux ad-ad is not torsion in $\mathcal{P}^{3}_{2,1}$, it is in special cases e.g. taking $a=d=e=f$. Although no analogous results have been brought to light tropically, classically [Beauville_torsion_Ceresa_class, Beauville_torsion_Ceresa_cycle] exhibit an example of a non-hyperelliptic curve which nevertheless has torsion Ceresa cycle.
@@
It is worth noting that a few years after Ceresa proved the inifinite order statement for the Ceresa cycle in $\text{Griff}_{1}(J(C))$, Bardelli proved that this group is in fact infinitely generated. He proceeds by constructing generators as follows: for every $k\geq 1$, the cycles $C_{k}-C_{k}^{-}\in\text{Griff}_{1}(J(C))$ are pushforwards of Ceresa cycles of curves $C_{k}$ for which $J(C_{k})$ has a $k$-to-11 map to $J(C)$. It is expected that one could prove a similar statement symplectically, only with our methods this would require a tropical construction of these infinitely many generators, which to the best of our knowledge has not appeared in the literature. There has been, however, progress on the tropical Schottky problem; for instance an algorithm to produce a tropical curve $C_{k}$ from the data of its Jacobian (and a theta divisor) is suggested in [bolognese2017curvestropicaljacobians, Section 7], and in [Chua2017SchottkyAC].